Rocky Corp. anticipates needing a new vehicle for its President 4 years from now. It believes the vehicle will cost $32,000 in four years. How much will the company need to save on a bi-annual (i.e., twice per year) basis starting six months from now if it believes it can earn 10 percent on the amount it saves? Ignore taxes.

$3,351 (rounded)

$6,895 (rounded)

$32,000 (rounded)

$545 (rounded)
Answer #1

Answer will be $3,351 (rounded)

Please find below table useful to compute desired results: -

Using PMT function of Excel

A BC 1 FV (Future Value) 32000 =-B1 2 NPER (number of periods) 4 =B2*2 3 RATE (Interest on savings) =10/100 =B3/2 4 PMT (mont

number of periods will be doubled and interest rate will be halved because it is a bi-annual basis

End results would be as follows: -

А BC 1 FV (Future Value) 32.000 (32,000) 2 NPER (number of periods) 3 RATE (Interest on savings) 10% 5% 4 PMT (monthly amount